If one spaceship fires a missile at 0.75c at the other spaceship as it approaches and the other spaceship observes the missile approaching it at 0.96c


Consider two spaceships moving relative to each other. The relative velocity of each spaceship is the velocity the other spaceship sees it moving at. Both relative velocities are equal in magnitude (equal speeds) but opposite in direction.


 If one spaceship fires a missile at 0.75c at the other spaceship as it approaches and the other spaceship observes the missile approaching it at 0.96c, what is the ratio of the first spaceship's relative velocity to the speed of light? Assume the velocity toward the observer is considered positive.
Answer #1

Relative velocity, u = (u' + v)/[1 + u'v/c^2]

u = (-0.75 c + 0.96 c)/[1 + (-0.75 x 0.96)c^2/c^2]

u = 0.21 c/0.28

u/c = 0.21/0.28

u/c = 0.75

Comment in case any doubt please rate my answer....